# Differentiable Cyclic Causal Discovery Under Unmeasured Confounders

### v1

@@ Abstract
Understanding causal relationships between variables is fundamental across scientific disciplines. Most causal discovery algorithms rely on two key assumptions: (i) all variables are observed, and (ii) the underlying causal graph is acyclic. While these assumptions simplify theoretical analysis, they are often violated in real-world systems, such as biological networks. Existing methods that account for confounders either assume linearity or struggle with scalability. To address these limitations, we propose DCCD-CONF, a novel framework for differentiable learning of nonlinear cyclic causal graphs in the presence of unmeasured confounders using interventional data. Our approach alternates between optimizing the graph structure and estimating the confounder distribution by maximizing the log-likelihood of the data. Through experiments on synthetic data and real-world gene perturbation datasets, we show that DCCD-CONF outperforms state-of-the-art methods in both causal graph recovery and confounder identification. Additionally, we also provide consistency guarantees for our framework, reinforcing its theoretical soundness.
@@ 1 Introduction
Modeling cause-effect relationships between variables is a fundamental problem in science [1, 2, 3], as it enables the prediction of a system’s behavior under previously unseen perturbations. These relationships are typically represented using directed graphs (DGs), where nodes correspond to variables, and directed edges capture causal dependencies. Consequently, causal discovery reduces to learning the structure of these graphs.
Existing causal discovery algorithms can be broadly classified into three categories: (i) constraint-based methods, (ii) score-based methods, and (iii) hybrid methods. Constraint-based methods, such as the PC algorithm [4, 5, 6], search for causal graphs that best satisfy the independence constraints observed in the data. However, since the number of conditional independence tests grows exponentially with the number of nodes, these methods often struggle with scalability. Score-based methods, such as the GES algorithm [7, 8], learn graph structures by maximizing a penalized score function, such as the Bayesian Information Criterion (BIC), over the space of graphs. Given the vast search space, these methods often employ greedy strategies to reduce computational complexity. A significant breakthrough came with Zheng et al. [9], who introduced a continuous constraint formulation to restrict the search space to acyclic graphs, inspiring several extensions [10, 11, 12, 13, 14, 15] that frame causal discovery as a continuous optimization problem under various model assumptions. Hybrid methods [16, 17, 18] integrate aspects of both constraint-based and score-based approaches, leveraging independence constraints while optimizing a score function.
@@
While these assumptions simplify the search space, they are often unrealistic, as real-world systems—especially in biology—frequently exhibit feedback loops and hidden confounders [19]. Enforcing these constraints can also increase computational complexity, particularly in ensuring acyclicity, which often requires solving challenging combinatorial or constrained optimization problems. These limitations hinder the practical applicability of existing methods in settings where such violations are unavoidable.
@@
Several approaches have been developed to address the challenge of feedback loops within causal graphs. Early work by Richardson [20] extended constraint-based approaches for Directed Acyclic Graphs (DAGs) to accommodate directed cycles. Another key contribution came from Lacerda et al. [21], who generalized Independent Component Analysis (ICA)-based causal discovery to handle linear non-Gaussian cyclic graphs. More recently, a growing body of research has focused on score-based methods for learning cyclic causal graphs [22, 23, 24, 25]. Additionally, some approaches leverage interventional data to improve structure recovery in cyclic systems. For instance, Hyttinen et al. [26] and Huetter and Rigollet [22] introduced frameworks that explicitly incorporate interventions to refine cyclic graph estimation. Sethuraman et al. [27] further advanced this line of research by introducing a differentiable framework for learning nonlinear cyclic graphs. Unlike differentiable DAG learners that enforce acyclicity through augmented Lagrangian-based solvers, their approach sidesteps these constraints by directly modeling the data likelihood, enabling more efficient and flexible learning of cyclic causal structures. However, their method assumes the absence of unmeasured confounders, which limits its applicability in real-world settings where hidden confounders are often present.
@@
Causal discovery in the presence of latent confounders has seen limited development, with most existing approaches grounded in constraint-based methodologies. Extensions of the PC algorithm, such as the Fast Causal Inference (FCI) algorithm [28], construct a Partial Ancestral Graph (PAG) to represent the equivalence class of DAGs in the presence of unmeasured confounders, and can accommodate nonlinear dependencies depending on the chosen conditional independence tests. However, standard FCI does not incorporate interventional data, prompting extensions such as JCI-FCI [29] and related approaches [30] that combine observational and interventional settings. Jaber et al. [31] further advanced this line of work by allowing for unknown interventional targets. Additionally, Forré and Mooij [32] introduced $\sigma$-separation, a generalization of $d$-separation, enabling constraint-based causal discovery in the presence of both cycles and latent confounders. A few recent approaches, such as Bhattacharya et al. [33], have explored continuous optimization frameworks using differentiable constraints, though these methods are currently limited to linear settings. Overall, a unified framework capable of handling nonlinearity, cycles, latent confounders, and interventions remains largely absent.
@@ 2.1 Structural Equations for Cyclic Causal Graphs
where $\text{pa}_{\mathcal{G}}(i):=\{j\in[d]:j\to i\in\mathcal{E}\}$ represents the parent set of $X_{i}$ in $\mathcal{G}$, and $\bm{X}_{\text{pa}_{\mathcal{G}}(i)}$ denotes the components of $\bm{X}=(X_{1},\ldots,X_{d})$ indexed by the parent set $\text{pa}_{\mathcal{G}}(i)$. We exclude self-loops (edges of the form $X_{i}\to X_{i}$) from $\mathcal{G}$, as their presence can lead to identifiability challenges [36]. The function $F_{i}$, referred to as causal mechanism, encodes the functional relationship between $X_{i}$ and its parents $\bm{X}_{\mathrm{pa}_{\mathcal{G}}(i)}$, and the exogenous noise variable $Z_{i}$.
The collection of exogenous noise variables $\bm{Z}=(Z_{1},\ldots,Z_{d})$ account for the stochastic nature as well as the confounding observed in the system. We make the assumption that the exogenous noise vector follows a Gaussian distributions: $\bm{Z}\sim\mathcal{N}(\bm{0},\mathbf{\Sigma}_{Z})$. Notably, if $(\mathbf{\Sigma}_{Z})_{ij}\neq 0$, then variables $X_{i}$ and $X_{j}$ are confounded, i.e., $i\leftrightarrow j\in\mathcal{B}$. In other words, confounding is modeled through correlations in the exogenous noise variables. Intuitively, if $X_{i}$ and $X_{j}$ share a hidden cause, their unexplained variation (the part not accounted for by their observed parents) will tend to move together. By allowing the noise terms $Z_{i}$ and $Z_{j}$ to be correlated, this shared influence can be effectively captured. This formulation generalizes prior work by allowing cycles, extending both nonlinear cyclic models without unmeasured confounders that assume independent noise terms [27], and acyclic models without confounders [37].
@@ 2.2 Interventions
In our work, we consider surgical interventions [35], also known as hard interventions, where all the incoming edges to the intervened nodes are removed from $\mathcal{G}$. Given a set of intervened upon nodes (also known as interventional targets), denoted as $I\subseteq\mathcal{V}$, the structural equations in (1) are modified as follows
@@
Assumption 2.1 (Interventional stability).
@@ 3.1 Modeling Causal Mechanism
We model the structural equations in (2) using implicit flows [38], which define an invertible mapping between $\bm{x}$ and $\bm{z}$ by solving the root of a function $\mathbf{G}(\bm{x},\bm{z})=\bm{0}$, where $\mathbf{G}:\mathbb{R}^{2d}\to\mathbb{R}^{d}$. Specifically, we take $\mathbf{G}(\bm{x},\bm{z})=\bm{x}-\mathbf{F}(\bm{x},\bm{z})$. General implicit mappings, however, do not guarantee invertibility or permit efficient computation of the log-determinant required for evaluating (6). To balance expressiveness with tractability, we adopt the structured form proposed by Lu et al. [38] for the causal mechanism:
@@
where $\mathbf{g}_{x}$ and $\mathbf{g}_{z}$ are restricted to be contractive functions. A function $\mathbf{g}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is contractive if there exists a constant $L<1$ such that $\|\mathbf{g}(\bm{x})-\mathbf{g}(\bm{y})\|\leq L\|\bm{x}-\bm{y}\|$ for all $\bm{x},\bm{y}\in\mathbb{R}^{d}$. This contractiveness ensures that the associated implicit map is uniquely solvable and invertible (see Theorem 1 in [38]).
Under this formulation, the forward map takes the form $\mathbf{f}_{x}(\bm{x})=(\mathbf{id}+\mathbf{g}_{z})^{-1}\circ(\mathbf{id}+\mathbf{g}_{x})(\bm{x})$, where $\mathbf{id}$ denotes the identity map. Given $\bm{x}$ (or $\bm{z}$), the corresponding value of $\bm{z}$ (or $\bm{x}$) can be computed via a root-finding procedure, i.e., $\bm{z}=\text{RootFind}(\bm{x}-\mathbf{F}(\bm{x},\cdot))$, specifically, we employ a quasi-Newton method (i.e., Broyden’s method [39]) to find the roots. To capture more complex nonlinear interactions between the observed variables $\bm{X}$ and latent confounders $\bm{Z}$, multiple such implicit blocks can be stacked. For simplicity, we focus on a single implicit flow block for subsequent discussion.
@@
where $\text{NN}(\cdot\mid\bm{\theta})$ denotes a fully connected neural network parameterized by $\bm{\theta}$, $\odot$ denotes the Hadamard product, and $\mathbf{M}^{\mathcal{G}}_{\ast,i}$ is the $i$-th column of $\mathbf{M}^{\mathcal{G}}$. The contractivity of $\mathbf{g}_{x}$ and $\mathbf{g}_{z}$ can be enforced by rescaling their weights using spectral normalization [40]. Moreover, the contractive nature of the causal mechanism facilitates efficient computation of the score function used for learning causal graphs, as discussed in Section 3.2.
@@ 3.2 Score function
Given a family of interventions $\mathcal{I}=\{I_{k}\}_{k\in[K]}$, we would like to learn the parameters of the structural equation model, i.e., causal graph structure, causal mechanism, and confounder distribution. To that end, similar to prior work [27, 41, 15] in this domain we employ regularized log-likelihood of the observed nodes as the score function to be maximized. That is,
@@
We now present the main theoretical result of this paper. The following theorem establishes that, under appropriate assumptions, the graph $\hat{\mathcal{G}}$ estimated by maximizing (9) belongs to the same general directed Markov equivalence class (introduced by [36]) as the ground truth graph $\mathcal{G}^{\ast}$ for each interventional setting $I_{k}\in\mathcal{I}$, denoted as $\hat{\mathcal{G}}\equiv_{\mathcal{I}}\mathcal{G}^{\ast}$, see Appendix A.1. Due to space constraints we provide the proof sketch below, see Appendix A.3 for complete proof of Theorem 3.1.
Theorem 3.1.
Let $\mathcal{I}=\{I_{k}\}_{k=1}^{K}$ be a family of interventional targets, let $\mathcal{G}^{\ast}$ denote the ground truth directed mixed graph, let $p^{(k)}$ denote the data generating distribution for $I_{k}$, and $\hat{\mathcal{G}}:=\arg\max_{\mathcal{G}}\mathcal{S}(\mathcal{G})$. Then, under the Assumptions 2.1, A.13, A.14, and A.15, and for a suitably chosen $\lambda>0$, we have that $\hat{\mathcal{G}}\equiv_{\mathcal{I}}\mathcal{G}^{\ast}$. That is, $\hat{\mathcal{G}}$ is $\mathcal{I}$-Markov equivalent to $\mathcal{G}^{\ast}$.
@@ 3.3 Updating model parameters
where we replace the expectation with respect to data distribution in (9) with sum over the finite samples, $\bm{x}^{(i,k)}$ represents the $i$-th data sample in the $k$-the interventional setting. We note that, since $p_{Z}=\mathcal{N}(\bm{0},\bm{\Sigma}_{Z})$, the covariance of the exogenous confounder vector, $\bm{\Sigma}_{Z}$, is implicitly embedded within $p_{\mathrm{do}(I_{k})\mathcal{G}}(\bm{x}^{(i,k)})$ in the score function.

### v2

@@ Abstract
Understanding causal relationships between variables is fundamental across scientific disciplines. Most causal discovery algorithms rely on two key assumptions: (i) all variables are observed, and (ii) the underlying causal graph is acyclic. While these assumptions simplify theoretical analysis, they are often violated in real-world systems, such as biological networks. Existing methods that account for confounders either assume linearity or struggle with scalability. To address these limitations, we propose DCCD-CONF, a novel framework for differentiable learning of nonlinear cyclic causal graphs in the presence of unmeasured confounders using interventional data. Our approach alternates between optimizing the graph structure and estimating the confounder distribution by maximizing the log-likelihood of the data. Through experiments on synthetic data and real-world gene perturbation datasets, we show that DCCD-CONF outperforms state-of-the-art methods in both causal graph recovery and confounder identification. Additionally, we provide consistency guarantees for our framework, reinforcing its theoretical soundness.
@@ 1 Introduction
Modeling cause-effect relationships between variables is a fundamental problem in science Sachs et al. [2005], Segal et al. [2005], Zhang et al. [2013], as it enables the prediction of a system’s behavior under previously unseen perturbations. These relationships are typically represented using directed graphs (DGs), where nodes correspond to variables, and directed edges capture causal dependencies. Consequently, causal discovery reduces to learning the structure of these graphs.
Existing causal discovery algorithms can be broadly classified into three categories: (i) constraint-based methods, (ii) score-based methods, and (iii) hybrid methods. Constraint-based methods, such as the PC algorithm [Spirtes et al., 2000, Triantafillou and Tsamardinos, 2015, Heinze-Deml et al., 2018], search for causal graphs that best satisfy the independence constraints observed in the data. However, since the number of conditional independence tests grows exponentially with the number of nodes, these methods often struggle with scalability. Score-based methods, such as the GES algorithm [Meek, 1997, Hauser and Bühlmann, 2012], learn graph structures by maximizing a penalized score function, such as the Bayesian Information Criterion (BIC), over the space of graphs. Given the vast search space, these methods often employ greedy strategies to reduce computational complexity. A significant breakthrough came with Zheng et al. [2018], who introduced a continuous constraint formulation to restrict the search space to acyclic graphs, inspiring several extensions [Yu et al., 2019, Ng et al., 2020, 2022, Zheng et al., 2020, Lee et al., 2019, Brouillard et al., 2020] that frame causal discovery as a continuous optimization problem under various model assumptions. Hybrid methods [Tsamardinos et al., 2006, Solus et al., 2017, Wang et al., 2017] integrate aspects of both constraint-based and score-based approaches, leveraging independence constraints while optimizing a score function.
@@
While these assumptions simplify the search space, they are often unrealistic, as real-world systems—especially in biology—frequently exhibit feedback loops and hidden confounders [Freimer et al., 2022]. Enforcing these constraints can also increase computational complexity, particularly in ensuring acyclicity, which often requires solving challenging combinatorial or constrained optimization problems. These limitations hinder the practical applicability of existing methods in settings where such violations are unavoidable.
@@
Several approaches have been developed to address the challenge of feedback loops within causal graphs. Early work by Richardson [1996] extended constraint-based approaches for Directed Acyclic Graphs (DAGs) to accommodate directed cycles. Another key contribution came from Lacerda et al. [2008], who generalized Independent Component Analysis (ICA)-based causal discovery to handle linear non-Gaussian cyclic graphs. More recently, a growing body of research has focused on score-based methods for learning cyclic causal graphs [Huetter and Rigollet, 2020, Améndola et al., 2020, Mooij and Heskes, 2013, Drton et al., 2019]. Additionally, some approaches leverage interventional data to improve structure recovery in cyclic systems. For instance, Hyttinen et al. [2012] and Huetter and Rigollet [2020] introduced frameworks that explicitly incorporate interventions to refine cyclic graph estimation. Sethuraman et al. [2023] further advanced this line of research by introducing a differentiable framework for learning nonlinear cyclic graphs. Unlike differentiable DAG learners that enforce acyclicity through augmented Lagrangian-based solvers, their approach sidesteps these constraints by directly modeling the data likelihood, enabling more efficient and flexible learning of cyclic causal structures. However, their method assumes the absence of unmeasured confounders, which limits its applicability in real-world settings where hidden confounders are often present.
@@
Causal discovery in the presence of latent confounders has seen limited development, with most existing approaches grounded in constraint-based methodologies. Extensions of the PC algorithm, such as the Fast Causal Inference (FCI) algorithm [Spirtes, 2001], construct a Partial Ancestral Graph (PAG) to represent the equivalence class of DAGs in the presence of unmeasured confounders, and can accommodate nonlinear dependencies depending on the chosen conditional independence tests. However, standard FCI does not incorporate interventional data, prompting extensions such as JCI-FCI [Mooij et al., 2020] and related approaches [Kocaoglu et al., 2019] that combine observational and interventional settings. Jaber et al. [2020] further advanced this line of work by allowing for unknown interventional targets. Additionally, Forré and Mooij [2018] introduced $\sigma$-separation, a generalization of $d$-separation, enabling constraint-based causal discovery in the presence of both cycles and latent confounders. Suzuki and Yang [2024] introduce LiNGAM-MMI, a generalization of the ICA-based LiNGAM Shimizu et al. [2006] that quantifies and mitigates confounding via a KL divergence minimization. Integer-programming–based formulations have also been proposed for causal discovery under latent confounding, such as [Ryšavỳ et al., 2025, Chen et al., 2021]. A few recent approaches, such as Bhattacharya et al. [2021], have explored continuous optimization frameworks using differentiable constraints, though these methods are currently limited to linear settings. In parallel, several works exist on causal inference under latent confounding—most notably Abadie et al. [2024], Chernozhukov et al. [2024]—propose doubly robust estimators that integrate outcome modeling, weighting, and cross-fitting for reliable effect estimation. Overall, a unified framework capable of handling nonlinearity, cycles, latent confounders, and interventions remains largely absent.
@@ 2.1 Structural Equations for Cyclic Causal Graphs
where $\text{pa}_{\mathcal{G}}(i):=\{j\in[d]:j\to i\in\mathcal{E}\}$ represents the parent set of $X_{i}$ in $\mathcal{G}$, and $\bm{X}_{\text{pa}_{\mathcal{G}}(i)}$ denotes the components of $\bm{X}=(X_{1},\ldots,X_{d})$ indexed by the parent set $\text{pa}_{\mathcal{G}}(i)$. We exclude self-loops (edges of the form $X_{i}\to X_{i}$) from $\mathcal{G}$, as their presence can lead to identifiability challenges [Bongers et al., 2021]. The function $F_{i}$, referred to as causal mechanism, encodes the functional relationship between $X_{i}$ and its parents $\bm{X}_{\mathrm{pa}_{\mathcal{G}}(i)}$, and the exogenous noise variable $Z_{i}$.
The collection of exogenous noise variables $\bm{Z}=(Z_{1},\ldots,Z_{d})$ account for the stochastic nature as well as the confounding observed in the system. We make the assumption that the exogenous noise vector follows a Gaussian distributions: $\bm{Z}\sim\mathcal{N}(\bm{0},\mathbf{\Sigma}_{Z})$. Notably, if $(\mathbf{\Sigma}_{Z})_{ij}\neq 0$, then variables $X_{i}$ and $X_{j}$ are confounded, i.e., $i\leftrightarrow j\in\mathcal{B}$. In other words, confounding is modeled through correlations in the exogenous noise variables. Intuitively, if $X_{i}$ and $X_{j}$ share a hidden cause, their unexplained variation (the part not accounted for by their observed parents) will tend to move together. By allowing the noise terms $Z_{i}$ and $Z_{j}$ to be correlated, this shared influence can be effectively captured. This formulation generalizes prior work by allowing cycles, extending both nonlinear cyclic models that assume independent noise terms [Sethuraman et al., 2023], and acyclic models without confounders [Peters and Bühlmann, 2014].
@@ 2.2 Interventions
In our work, we consider surgical interventions [Pearl, 2009], also known as hard interventions, where all the incoming edges to the intervened nodes are removed from $\mathcal{G}$. Given a set of intervened upon nodes (also known as interventional targets), denoted as $I\subseteq\mathcal{V}$, the structural equations in (1) are modified as follows
@@
Assumption 1 (Interventional stability).
@@ 3.1 Modeling Causal Mechanism
We model the structural equations in (2) using implicit flows [Lu et al., 2021], which define an invertible mapping between $\bm{x}$ and $\bm{z}$ by solving the root of a function $\mathbf{G}(\bm{x},\bm{z})=\bm{0}$, where $\mathbf{G}:\mathbb{R}^{2d}\to\mathbb{R}^{d}$. Specifically, we take $\mathbf{G}(\bm{x},\bm{z})=\bm{x}-\mathbf{F}(\bm{x},\bm{z})$. General implicit mappings, however, do not guarantee invertibility or permit efficient computation of the log-determinant required for evaluating (6). To balance expressiveness with tractability, we adopt the structured form proposed by Lu et al. [2021] for the causal mechanism:
@@
where $\mathbf{g}_{x}$ and $\mathbf{g}_{z}$ are restricted to be contractive functions. A function $\mathbf{g}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is contractive if there exists a constant $L<1$ such that $\|\mathbf{g}(\bm{x})-\mathbf{g}(\bm{y})\|\leq L\|\bm{x}-\bm{y}\|$ for all $\bm{x},\bm{y}\in\mathbb{R}^{d}$. This contractiveness ensures that the associated implicit map is uniquely solvable and invertible (see Theorem 1 in [Lu et al., 2021]). In other words, contractivity ensures that the process defined by the SEM converges to an equilibrium state.
Under this formulation, the forward map takes the form $\mathbf{f}_{x}(\bm{x})=(\mathbf{id}+\mathbf{g}_{z})^{-1}\circ(\mathbf{id}+\mathbf{g}_{x})(\bm{x})$, where $\mathbf{id}$ denotes the identity map. Given $\bm{x}$ (or $\bm{z}$), the corresponding value of $\bm{z}$ (or $\bm{x}$) can be computed via a root-finding procedure, i.e., $\bm{z}=\text{RootFind}(\bm{x}-\mathbf{F}(\bm{x},\cdot))$, specifically, we employ a quasi-Newton method (i.e., Broyden’s method [Broyden, 1965]) to find the root. To capture more complex nonlinear interactions between the observed variables $\bm{X}$ and latent confounders $\bm{Z}$, multiple such implicit blocks can be stacked. This is true since $\mathbf{f}_{x}$ is highly nonlinear and by suitably parameterizing $\mathbf{g}_{x}$ and $\mathbf{g}_{z}$ any nonlinear interaction between $\bm{x}$ and $\bm{z}$ can be modeled. For simplicity, we focus on a single implicit flow block for subsequent discussion.
@@
where $\text{NN}(\cdot\mid\bm{\theta})$ denotes a fully connected neural network parameterized by $\bm{\theta}$, $\odot$ denotes the Hadamard product, and $\mathbf{M}^{\mathcal{G}}_{\ast,i}$ is the $i$-th column of $\mathbf{M}^{\mathcal{G}}$. The contractivity of $\mathbf{g}_{x}$ and $\mathbf{g}_{z}$ can be enforced by rescaling their weights using spectral normalization [Behrmann et al., 2019]. Moreover, the contractive nature of the causal mechanism facilitates efficient computation of the score function used for learning causal graphs, as discussed in Section 3.2.
@@ 3.2 Score function
Given a family of interventions $\mathcal{I}=\{I_{k}\}_{k\in[K]}$, we would like to learn the parameters of the structural equation model, i.e., causal graph structure, causal mechanism, and confounder distribution. To that end, similar to prior work [Sethuraman et al., 2023, Lopez et al., 2022, Brouillard et al., 2020] in this domain we employ regularized log-likelihood of the observed nodes as the score function to be maximized. That is,
@@
We now present the main theoretical result of this paper. The following theorem establishes that, under appropriate assumptions, the graph $\hat{\mathcal{G}}$ estimated by maximizing (9) belongs to the same general directed Markov equivalence class (introduced by [Bongers et al., 2021]) as the ground truth graph $\mathcal{G}^{\ast}$ for each interventional setting $I_{k}\in\mathcal{I}$, denoted as $\hat{\mathcal{G}}\equiv_{\mathcal{I}}\mathcal{G}^{\ast}$, see Appendix A.1. Due to space constraints we provide the proof sketch below, see Appendix A.3 for complete proof of Theorem 2.
Theorem 2.
Let $\mathcal{I}=\{I_{k}\}_{k=1}^{K}$ be a family of interventional targets, let $\mathcal{G}^{\ast}$ denote the ground truth directed mixed graph, let $p^{(k)}$ denote the data generating distribution for $I_{k}$, and $\hat{\mathcal{G}}:=\arg\max_{\mathcal{G}}\mathcal{S}(\mathcal{G})$. Then, under the Assumptions 1, A.13, A.14, and A.15, and for a suitably chosen $\lambda>0$, we have that $\hat{\mathcal{G}}\equiv_{\mathcal{I}}\mathcal{G}^{\ast}$. That is, $\hat{\mathcal{G}}$ is $\mathcal{I}$-Markov equivalent to $\mathcal{G}^{\ast}$.
Theorem 2 rests on three key assumptions. Assumption A.13 ensures that the data-generating distribution lies within the model class, while Assumption A.14 guarantees that every statistical independence in the data corresponds to a $\sigma$-separation in the ground-truth graph. Finally, Assumption 1 prevents the score function from diverging to infinity.
@@ 3.3 Updating model parameters
where we replace the expectation with respect to data distribution in (9) with sum over the finite samples, $\bm{x}^{(i,k)}$ represents the $i$-th data sample in the $k$-the interventional setting. We note that, since $p_{Z}=\mathcal{N}(\bm{0},\bm{\Sigma}_{Z})$, the covariance of the exogenous confounder vector, $\bm{\Sigma}_{Z}$, is implicitly embedded within $p_{\mathrm{do}(I_{k})(\mathcal{G})}(\bm{x}^{(i,k)})$ in the score function.